# Solvation Free Energies from Neural Thermodynamic Integration

### v1

@@ Abstract
We propose to compute solvation free energies via thermodynamic integration along a neural-network potential interpolating between two target Hamiltonians. We use a stochastic interpolant to define an interpolation between the distributions at the level of samples and optimize a neural network potential to match the corresponding equilibrium potential at every intermediate time-step. Once the alignment between the interpolating samples and the interpolating potentials is sufficiently accurate, the free-energy difference between the two Hamiltonians can be estimated using (neural) thermodynamic integration. We validate our method to compute solvation free energies on several benchmark systems: a Lennard-Jones particle in a Lennard-Jones fluid, as well as the insertion of both water and methane solutes in a water solvent at atomistic resolution.
1 Introduction
Estimating free energy differences is at the heart of understanding a wide range of physical, chemical, and biological processes, from protein folding and ligand binding to phase transitions in materials. These calculations offer invaluable insights into the stability of molecular conformations, the spontaneity of chemical reactions, and the mechanisms that drive phase changes (Gao et al., 2006; Mobley & Gilson, 2017; Agarwal et al., 2021). The traditional approaches for performing such calculations, such as thermodynamic integration (TI) and free energy perturbation (FEP) have recently been complemented with ideas from the machine learning community (Noé et al., 2019; Wirnsberger et al., 2020; Invernizzi et al., 2022).
@@ 1 Introduction
This work generalizes NeuralTI (Máté et al., 2024), a machine-learning enhanced variant of TI, to estimate the free energy difference between two arbitrary Hamiltonians. In §3 we argue that stochastic interpolants provide a natural framework for this setting and describe our approach. Finally, in §4 we validate our methodology by computing solvation free energies on several systems: ($i$) a Lennard-Jones (LJ) solute in an LJ solvent, and the insertion of ($ii$) a water solute and ($iii$) a methane solute in a box of water, both at atomistic resolution. All free energies show excellent agreement with reference calculations.
@@
Figure 1: An interpolating family of distributions coupling a solute (brown) to the solvent (blue). The potentials $U_{A},U_{B}$ and $U_{AB}$ denote the interactions within the solvent, within the solute and between the two components, respectively. The interpolation $U_{t}=U_{A}+U_{B}+tU_{AB}$ is often used in TI calculations to compute the free energy of the coupling of the solute, we include an additional trainable potential $t(1-t)U_{t}^{\theta}$ and train it to be the equilibrium potential of a stochastic interpolant.
2 Background
@@
We begin with fixing some notation. Depending on which arguments are important in the context, we will use $U_{t}(x)$ or $U(t,x)$ or simply $U$ to denote time-dependent energy/potential functions and $\rho_{t}$ for the corresponding Boltzmann probability density $\frac{1}{Z_{t}}e^{-\beta U_{t}}$, with inverse temperature $\beta=1/k_{B}T$. The symbol $\nabla U$ is used for the spatial gradient. Isotropic Gaussian distributions with variance $\sigma^{2}$ are written as $\mathcal{G}_{\sigma}$ and the convolution of two functions, $f$ and $g$ as $f\star g$.
@@
Thermodynamic integration (Kirkwood, 1935) is a method for estimating the free energy difference between two Hamiltonians relying on a family of potentials, $U_{\lambda}$, interpolating between $U_{0}$ and $U_{1}$.
@@
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expected value of $\partial_{\lambda}U_{\lambda}$ under $\rho_{\lambda}=Z_{\lambda}^{-1}e^{-\beta U_{\lambda}}$. To numerically estimate the value of the last integral one needs an interpolating potential function $U_{\lambda}$ and samples from the corresponding interpolating equilibrium distributions $\rho_{\lambda}$. Usually one first fixes an interpolation $U_{\lambda}$ and performs Monte Carlo sampling or molecular dynamics (MD) simulations at intermediate $\lambda$ values to end up with the estimate
$\Delta\hat{F}_{0\rightarrow 1}=\frac{1}{N+1}\sum_{\lambda_{i}}\,\mathbb{E}_{x\sim\rho_{\lambda_{i}}}\big{[}\partial_{\lambda}U_{\lambda}(x)\big{]}\qquad\lambda_{i}\in\{0,1/N,2/N,...,1\}$ (3)
@@
Máté et al. (2024) proposed to parametrize a time-dependent energy function $U_{t}$ and use denoising score matching to align the force $\nabla U_{t}$ with the scores of increasingly noised versions of the target Boltzmann distribution $e^{-\beta U_{0}}/Z_{0}$. As the noise level $\sigma(t)$ starts small at $t=0$, and is large at $t=1$, the trained $U_{t}$ provides and interpolating potential between the target and a tractable prior. Since samples from the corresponding Boltzmann distributions $e^{-\beta U_{t}}\approx\mathcal{G}_{\sigma_{t}}\star e^{-\beta U_{0}}$ are also (approximately) available, TI can be performed along $U_{t}$ to estimate the free-energy difference between a trivial and a non-trivial distribution. In this case, these were matched to the ideal gas and the Boltzmann distribution of the fully coupled LJ liquids, so as to associate the resulting free energy to the coupling of all the interactions.
Traditional approaches to solvation free-energy calculations often consist of first fixing the interpolation in the space of potentials (e.g., linear coupling), to subsequently generate samples from Monte Carlo sampling or molecular dynamics simulations Mey et al. (2020). Here instead, we first fix the sampling procedure and learn the interpolating family of potentials.
3 Method
@@
The goal of this work is to extend the neural TI framework to the estimation of free energy differences between two non-trivial Hamiltonians. To this end, we consider a pair of potential functions, $U_{0},U_{1}:\mathcal{M}\rightarrow\mathbb{R}$ on a manifold $\mathcal{M}$ representing the configuration space of some system. In our experiments, we model particle systems and rigid molecules in a periodic box. For the particle system $\mathcal{M}$ will be the $3N$-torus, $\mathbb{T}^{3N}$ representing the positions of $N$ particles in a 3-torus. For the rigid molecular system, the configuration of each molecule is given by its position in $\mathbb{T}^{3}$ and its orientation, thus the configuration space of the full system is $\mathcal{M}=[\mathbb{T}^{3}\times SO(3)]^{N}$.
Stochastic Interpolants
We assume that we can sample $x_{0}\sim\rho_{0}\propto e^{-\beta U_{0}}$ and $x_{1}\sim\rho_{1}\propto e^{-\beta U_{1}}$. We then define an interpolating family of distributions $\rho_{t}$ for $10<t<$ 1 by constructing a procedure to sample $x_{t}$ using a stochastic interpolant (Albergo & Vanden-Eijnden, 2023; Albergo et al., 2023). For the sake of simplicity, let us first assume that $\mathcal{M}=\mathbb{R}^{n}$. A stochastic interpolant generates interpolating samples from $\rho_{t}$ by
$x_{t}=I(x_{0},x_{1},t)+\sigma_{t}z$ (4)
@@
where $x_{0}\sim\rho_{0},x_{1}\sim\rho_{1},z\sim\mathcal{N}(0,1)$, $I$ satisfies $I(x_{0},x_{1},0)=x_{0},I(x_{0},x_{1},1)=x_{1}$, e.g., $I(x_{0},x_{1},t)=(1-t)x_{0}+tx_{1}$ and $\sigma_{0},\sigma_{1}\ll 1$. If $\sigma_{0}=\sigma_{1}=0$, then the boundary conditions are exactly satisfied, but $0<\sigma_{0},\sigma_{1}\ll 1$ also allows score-based training, while only slightly violating the boundary conditions. Conditioned on $x_{0}$ and $x_{1}$, $\rho_{t}(x_{t}|x_{0},x_{1})$ is an isotropic Gaussian with mean $I(x_{0},x_{t},t)$, variance $\sigma_{t}$ and Stein-score $\nabla\log\rho_{t}(x_{t}|x_{0},x_{1})=\frac{1}{\sigma_{t}}(I(x_{0},x_{1},t)-x_{t})=-z/\sigma_{t}$. The marginal score $\nabla\log\rho_{t}(x)$ reads
$\nabla\log\rho_{t}(x)=\mathbb{E}_{x_{0},x_{1}}\left[\nabla\log\rho_{t}(x_{t}|x_{0},x_{1})\right]=\mathbb{E}_{x_{0},x_{1}}\left[-z/\sigma_{t}\right],$ (5)
@@
enabling the training of an energy-based model using the usual denoising score matching (DSM) objective (Vincent, 2011). This approach can be naturally extended to non-Euclidean $\mathcal{M}$ by working in local charts on the manifold (De Bortoli et al., 2022).
@@
The DSM objective requires the computation of the spatial gradient $\nabla_{x}U_{t}(x)$ at samples $x_{t}$, sampled at $(t,x)\sim(\mathcal{U}([0,1]),\rho_{t})$. The temporal derivative of the energy function, $\partial_{t}U_{t}(x)$, can be easily calculated thanks to automatic differentiation at the same points, yielding a Monte Carlo estimate of the TI integral in Equation (2),
$\Delta\hat{F}_{0\rightarrow 1}=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x\sim\rho_{t}}\big{[}\partial_{t}U_{t}(x)\big{]}$ (6)
In our experiments, at every training step we compute both the spatial and temporal derivatives of $U_{t}$ and use $\nabla U_{t}$ for the DSM objective and $\partial_{t}U_{t}$ for a rolling estimate of the free-energy difference.
Locally consistent potentials
Our strategy relies on learning a time-dependent potential $U_{t}$ corresponding to the samples $x_{t}\sim\rho_{t}$ produced by a stochastic interpolant. To perform TI, the samples $x_{t}$ and the energy function $\beta U_{t}$ need to be compatible with each other in the sense that $x_{t}\sim e^{-\beta U_{t}}$. An energy-based model trained with the DSM objective can only fit an energy function up to an additive constant. Since we aim to learn a time-dependent family of energies, the degrees of freedom correspond to a time-dependent function defining the energy scale along the trajectory. Because we enforce the boundary conditions by construction (equation 12), this degree of freedom vanishes when integrated over the full time interval $[0,1]$.
A possible issue arises for a target domain that is disconnected or a distribution that is multimodal: the additive constant becomes local. Thus, strictly speaking, our model does not learn a function $U_{t}$ such that $e^{-\beta U_{t}}$ is globally proportional to $\rho_{t}$, but only locally proportional. In other words, it does not preserve the relative weights of the modes, but does preserve the relative likelihoods within the modes. (Song & Ermon, 2019) solve this by learning a family of models with increasing noise levels, making use of the fact that for high perturbations the target distribution becomes connected. Alternatively, Gutmann & Hyvärinen (2010) overcome the issue by contrasting the data with a second distribution of known density. In the present work, the enforced boundary conditions alleviate the issue: All offsets introduced to the relative weighting of the modes also integrate to zero over the time interval $[0,1]$. See Figure 2 for a one-dimensional illustration of this argument.
Figure 2: Unnormalized interpolating densities $e^{-\beta U_{t}^{\theta}}$ learned from different initializations. The two target potentials are given by a standard Gaussian at $t=0$ ($\log Z_{0}=1$) and the sum of two Gaussian densities at $t=1$ ($\log Z_{1}=2$). The 4 different models learn 4 different relative weightings of the modes along the interpolation, but have a consistent prediction for the free energy difference of $\log 2\approx 0.69$.
4 Experiments
@@
$U_{1}$ $=U_{\mathrm{solvent}}+U_{\mathrm{solute}}+U_{\mathrm{solute-solvent}}$ (8)
where $U_{\mathrm{solvent}},U_{\mathrm{solute}}$ and $U_{\mathrm{solute-solvent}}$ represent the interactions within the solvent, within the solute and cross interactions between the two components, respectively.
4.1 Solute solvation of a Lennard-Jones fluid
System description
We consider a system of 512 solvent particles and an additional solute particle with Lennard-Jones (LJ) interactions between them. The LJ parameters of the solvent are denoted by $(\varepsilon_{A},\sigma_{A})$ while for the solute $(\varepsilon_{B},\sigma_{B})=(2\varepsilon_{A},2\sigma_{A})$ and the LJ interactions are computed by the Lorentz-Berthelot rules (Lorentz, 1881; Berthelot, 1898) $(\varepsilon_{AB}=\sqrt{\varepsilon_{A}\varepsilon_{B}},\sigma_{AB}=(\sigma_{A}+\sigma_{B})/2)$. The simulation box is of size $9\sigma_{A}\times 9\sigma_{A}\times 9\sigma_{A}$ with periodic boundary conditions. The potential functions $U_{\mathrm{solvent}},U_{\mathrm{solute}}+U_{\mathrm{solute-solvent}}$ are then given by
$U_{\mathrm{solute}}$ $=0$ (9)
$U_{\mathrm{solvent}}$ $=\sum_{i\neq j}4\varepsilon_{A}\left[\left(\frac{\sigma_{A}}{r_{ij}}\right)^{12}-\left(\frac{\sigma_{A}}{r_{ij}}\right)^{6}\right]$ (10)
$U_{\mathrm{solute-solvent}}$ $=\sum_{i}4\varepsilon_{AB}\left[\left(\frac{\sigma_{AB}}{r_{i}}\right)^{12}-\left(\frac{\sigma_{AB}}{r_{i}}\right)^{6}\right]$ (11)
where $r_{ij}$ is the distance between the solvent particles with indices $i$ and $j$, and $r_{i}$ is the distance between the solute and the solvent particle with index $i$.

### v2

@@ Abstract
We present a method for computing free-energy differences using thermodynamic integration with a neural network potential that interpolates between two target Hamiltonians. The interpolation is defined at the sample distribution level, and the neural network potential is optimized to match the corresponding equilibrium potential at every intermediate time-step. Once the interpolating potentials and samples are well-aligned, the free-energy difference can be estimated using (neural) thermodynamic integration. To target molecular systems, we simultaneously couple Lennard-Jones and electrostatic interactions and model the rigid-body rotation of molecules. We report accurate results for several benchmark systems: a Lennard-Jones particle in a Lennard-Jones fluid, as well as the insertion of both water and methane solutes in a water solvent at atomistic resolution using a simple three-body neural-network potential.
preprint: AIP/123-QED**footnotetext: balint.mate@unige.chfootnotetext: francois.fleuret@unige.chfootnotetext: bereau@uni-heidelberg.de
I Introduction
Estimating free-energy differences is at the heart of understanding a wide range of physical, chemical, and biological processes, from protein folding and ligand binding to phase transitions in materials. These calculations offer invaluable insights into the stability of molecular conformations, the spontaneity of chemical reactions, and the mechanisms that drive phase changes [1, 2, 3, 4]. The naïve way to compute free-energy differences would be to subtract two individual free energies. The issue with this approach is that the individual free energies are usually orders-of-magnitude larger than their difference. While theoretically sound, it requires extreme precision to predict a small quantity as the difference of two larger ones. Instead a variety of methods have been developed to compute free-energy differences directly. For instance, transfer free energies between different fluids [5, 6], solvation or hydration free energies [7, 8, 9, 10], or protein-ligand binding [11, 12, 13, 14]. Estimating free energies or free-energy differences involves integrals over large configuration spaces and is usually evaluated using techniques from statistical physics such as thermodynamic integration (TI) and free-energy perturbation (FEP) [15]. These computations boil down to the task of sampling from equilibrium distributions and using the samples to perform numerical integration. To bypass the computation of such configurational integrals, it has been proposed to train machine-learning (ML) models on experimental or simulation labels to directly regress free energies [16, 17, 18, 19, 20]. Another promising class of approaches seeks to replace classical sampling methods, such as Markov Chain Monte Carlo (MCMC) and molecular dynamics (MD), with techniques inspired by the machine learning community [21, 22, 23]. These methods leverage generative models to produce i.i.d samples from their target distributions, enabling more effective computation of configurational integrals.
We build on NeuralTI [24], an approach performing TI using a neural-network potential, to estimate the free-energy difference between two arbitrary Hamiltonians. Thermodynamic integration estimates a free-energy difference between two Hamiltonians, $\mathcal{H}_{0},\mathcal{H}_{1}$, defined on the same configuration space, via a one-parameter family of Hamiltonians $\mathcal{H}_{\lambda}$ interpolating between $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$. Given such an interpolation, the free-energy difference can then be written as an integral along the coupling interval $\Delta F_{0\to 1}=\int_{0}^{1}\mathrm{d}\lambda\langle\partial\mathcal{H}_{\lambda}/\partial\lambda\rangle_{\lambda}$, where the brackets .$\langle.\rangle_{\lambda}$ denote the expectation value with respect to the Boltzmann distribution $e^{-\beta\mathcal{H}_{\lambda}}/{Z_{\lambda}}$. To numerically evaluate the TI integral, one usually discretizes the coupling variable and generates samples from the intermediate Boltzmann distributions at various values of $\lambda$. The idea behind NeuralTI is to progressively transform samples between the endpoint distributions, learn the corresponding equilibrium $\mathcal{H}_{\lambda}$ with a neural network, and perform TI using the learnt $\mathcal{H}_{\lambda}$. This idea was first applied to a denoising diffusion model interpolating between the ideal gas and a Lennard-Jones (LJ) fluid in a periodic box [24]. This approach relied on the observation that the Hamiltonian of the ideal gas is trivial thus the corresponding Boltzmann distribution could be easily identified with the latent space of a diffusion model.
@@ I Introduction
In this work, we apply the same approach to a scenario where both endpoint Hamiltonians are not trivial, but represent coupled and decoupled states of a solute–solvent system as illustrated in Fig. 1. We generate samples $x_{t}\sim\rho_{t}$ as the geodesic interpolation between samples $x_{0}\sim\rho_{0},x_{1}\sim\rho_{1}$ and optimize a neural network potential to match the corresponding equilibrium potential at every $10<t<$ 1. Notably, in our experiments on chemical systems, the Lennard-Jones and electrostatic interactions are coupled simultaneously as parametrized by the neural network. The paper is structured as follows: Section II summarizes the relevant concepts that our work builds on. In Section III we argue that ideas from the machine-learning community generalizing diffusion models provide a natural framework for this setting [25, 26, 27] and describe our approach. Finally, in Section IV we validate our methodology by computing solvation free energies on several systems: ($i$) an LJ solute in an LJ solvent, and the insertion of ($ii$) a water solute and ($iii$) a methane solute in a box of water, both at atomistic resolution. All free energies show excellent agreement with reference free energies.
@@
Figure 1: An interpolating family of distributions coupling a solute (brown) to the solvent (blue). The potentials $U_{A},U_{B}$ and $U_{AB}$ denote the interactions within the solvent, within the solute and between the two components, respectively. The interpolation $U_{t}=U_{A}+U_{B}+tU_{AB}$ is often used in TI calculations to compute the free energy of the coupling of the solute, we include an additional trainable potential $t(1-t)U_{t}^{\theta}$ and train it to be the equilibrium potential at all intermediate time-slices.
II Background
@@
We begin with fixing some notation. Depending on which arguments are important in the context, we will use $U_{t}(x)$ or $U(t,x)$ or simply $U$ to denote time-dependent energy/potential functions and $\rho_{t}$ for the corresponding Boltzmann probability density $\frac{1}{Z_{t}}e^{-\beta U_{t}}$, with inverse temperature $\beta=1/k_{\mathrm{B}}T$. The symbol $\nabla U$ is used for the spatial gradient. The isotropic Gaussian distribution, centered at the origin, with variance $\sigma^{2}$, is written as $\mathcal{N}_{\sigma^{2}}$ and the convolution of two functions, $f$ and $g$ as $f\star g$.
@@
Thermodynamic integration [28] is a method for estimating the free-energy difference between two Hamiltonians relying on a family of potentials, $U_{\lambda}$, interpolating between $U_{0}$ and $U_{1}$.
@@
where $\langle\partial_{\lambda}U_{\lambda}\rangle_{\lambda}$ denotes the expected value of $\partial_{\lambda}U_{\lambda}$ under $\rho_{\lambda}=Z_{\lambda}^{-1}e^{-\beta U_{\lambda}}$. To numerically estimate the value of the last integral one needs an interpolating potential function $U_{\lambda}$ and samples from the corresponding interpolating equilibrium distributions $\rho_{\lambda}$. Usually one first fixes an interpolation $U_{\lambda}$ and performs Monte Carlo sampling or molecular dynamics (MD) simulations at intermediate $\lambda$ values to obtain the estimate
$\Delta\hat{F}_{0\rightarrow 1}=\frac{1}{N+1}\sum_{\lambda_{i}}\,\mathbb{E}_{x\sim\rho_{\lambda_{i}}}\big{[}\partial_{\lambda}U_{\lambda}(x)\big{]},$ (3)
where the sum is over $\lambda_{i}\in\{0,1/N,2/N,...,1\}$.
@@
Máté et al. [24] proposed to parametrize a time-dependent energy function $U_{t}$ and use score matching (SM) to align the force $\nabla U_{t}$ with the scores of increasingly noised versions of the target Boltzmann distribution $e^{-\beta U_{0}}/Z_{0}$. As the noise level $\sigma(t)$ starts small at $t=0$, and is large at $t=1$, the trained $U_{t}$ provides an interpolating potential between the target and a tractable prior. Since samples from the corresponding Boltzmann distributions $e^{-\beta U_{t}}\approx\mathcal{N}_{\sigma_{t}^{2}}\star e^{-\beta U_{0}}$ are also (approximately) available, TI can be performed using $U_{t}$ to estimate the free-energy difference between a trivial and a non-trivial distribution. In this case, these were matched to the ideal gas and the Boltzmann distribution of the fully coupled LJ liquids, so as to associate the resulting free energy to the coupling of all the interactions. Traditional approaches to solvation free-energy calculations often consist of first fixing the interpolation in the space of potentials (e.g., linear coupling), to subsequently generate samples from Monte Carlo sampling or molecular dynamics simulations [15]. Here instead, we first fix the sampling procedure and learn the interpolating family of potentials.
Locally consistent potentials
The above strategy relies on learning a time-dependent potential $U_{t}$ corresponding to samples $x_{t}\sim\rho_{t}$. To perform TI, the samples $x_{t}$ and the energy function $\beta U_{t}$ need to be compatible with each other in the sense that $x_{t}\sim\frac{1}{Z_{t}}e^{-\beta U_{t}}$. An energy-based model trained with an SM objective can only fit an energy function up to an additive constant. Since we aim to learn a time-dependent family of energies, the degrees of freedom correspond to a time-dependent function defining the energy scale along the trajectory. Because the boundary conditions are enforced by construction, the integral of this degree of freedom over the full time interval $[0,1]$ must vanish.
A possible issue arises for a domain that is disconnected or a target distribution that is multimodal: the additive constant becomes local. Thus, strictly speaking, our model does not learn a function $U_{t}$ such that $e^{-\beta U_{t}}$ is globally proportional to $\rho_{t}$, but only locally proportional. In other words, it does not preserve the relative weights of the modes, but does preserve the relative likelihoods within the modes. Song and Ermon [29] solve this by learning a family of models with increasing noise levels, making use of the fact that for high perturbations the target distribution becomes connected. Alternatively, Gutmann and Hyvärinen [30] overcome the issue by contrasting the data with a second distribution of known density. In the present work, the enforced boundary conditions alleviate the issue: All offsets introduced to the relative weighting of the modes also integrate to zero over the time interval $[0,1]$. See Figure 2 for a one-dimensional illustration.
Figure 2: Unnormalized interpolating densities $e^{-\beta U_{t}^{\theta}}$ learned from different initializations. The two endpoint potentials are given by a standard Gaussian at $t=0$ ($\log Z_{0}=1$) and the sum of two Gaussian densities at $t=1$ ($\log Z_{1}=2$). The 4 different models learn 4 different relative weightings of the modes along the interpolation, but have a consistent prediction for the free-energy difference of $\log 2\approx 0.69$.
III Method
@@
The goal of this work is to extend the neural TI framework to the estimation of free-energy differences between two non-trivial Hamiltonians. To this end, we consider a pair of potential functions, $U_{0},U_{1}:\mathcal{M}\rightarrow\mathbb{R}$ on a manifold $\mathcal{M}$ representing the configuration space of some system. In our experiments, we model particle systems and rigid molecules in a periodic box. For the particle system $\mathcal{M}$ will be the $3N$-torus, $\mathbb{T}^{3N}$ representing the positions of $N$ particles in a 3-torus. For the rigid molecular system, the configuration of each molecule is given by its position in $\mathbb{T}^{3}$ and its orientation, thus the configuration space of the full system is $\mathcal{M}=[\mathbb{T}^{3}\times SO(3)]^{N}$.
Learning an interpolating potential
Sampling the interpolating densities
Given two densities $\rho_{0},\rho_{1}$, there are infinitely many possible interpolations between them. To obtain a unique and well-defined interpolation, one could impose additional conditions on the transport [31, 32]. However, since this work focuses on computing the free-energy difference between the endpoints, all interpolations are equally valid, as the free energy is a state function. Consequently, our design choices regarding the interpolation are driven solely by considerations of computational simplicity. For what follows, we assume that we have access to samples $x_{0}\sim\rho_{0}\propto e^{-\beta U_{0}}$ and $x_{1}\sim\rho_{1}\propto e^{-\beta U_{1}}$. We then define an interpolating family of distributions $\rho_{t}$ for $10<t<$ 1 by constructing a procedure to sample $x_{t}$ a linear interpolation between a pair of samples from the endpoint distributions [26, 27, 25]. We then construct interpolating samples by
$x_{t}=I(x_{0},x_{1},t),$ (4)
where $x_{0}\sim\rho_{0},x_{1}\sim\rho_{1}$, and $I$ satisfies $I(x_{0},x_{1},0)=x_{0},I(x_{0},x_{1},1)=x_{1}$, e.g., $I(x_{0},x_{1},t)=(1-t)x_{0}+tx_{1}$ on $\mathbb{R}^{n}$ and the geodesic interpolation on $\mathbb{T}^{3}$ and $SO(3)$. To learn the corresponding equilibrium potentials we employ score matching (SM).
Denoising Score matching
Denoising score matching (DSM) allows us to learn the score of a noised version of $\rho_{t}$
$\tilde{\rho}_{t}=\rho_{t}\star\mathcal{N}_{\sigma_{t}^{2}}.$ (5)
@@
Performing TI along $\tilde{\rho}_{t}$ approximates TI along $\rho_{t}$ as long as $\sigma_{0}\approx 0$ and $\sigma_{1}\approx 0$. If $\sigma_{0}=\sigma_{1}=0$, then the boundary conditions are exactly satisfied, but DSM requires a non-zero $\sigma$, thus we choose $0<\sigma_{0},\sigma_{1}\ll 1$ only slightly violating the boundary conditions. Conditioned on $x_{0}$ and $x_{1}$, $\tilde{\rho}_{t}(x_{t}|x_{0},x_{1})$ is an isotropic Gaussian with mean $I(x_{0},x_{t},t)$, variance $\sigma_{t}$ and Stein-score $\nabla\log\tilde{\rho}_{t}(x_{t}|x_{0},x_{1})=\frac{1}{\sigma_{t}}(I(x_{0},x_{1},t)-x_{t})=-z/\sigma_{t}$. The marginal score $\nabla\log\tilde{\rho}_{t}(x)$ reads
$\nabla\log\tilde{\rho}_{t}(x)=\mathbb{E}_{x_{0},x_{1}}\left[\nabla\log\tilde{\rho}_{t}(x_{t}|x_{0},x_{1})\right]=\mathbb{E}_{x_{0},x_{1}}\left[-z/\sigma_{t}\right],$ (6)
@@
enabling the training of an energy-based model using the usual DSM objective [33]. This approach can be naturally extended to non-Euclidean $\mathcal{M}$ by learning the score in local charts on the manifold [34].
Target Score matching
De Bortoli et al. [35] recently proposed target score matching (TSM), an alternative SM objective that is well-suited for our application. TSM relies on having access to target score functions at the endpoints, $\nabla\log\rho_{i}=-\nabla U_{i}$ for $i\in\{0,1\}$. Decomposing $\nabla\log\rho_{t}$, as
$\nabla\log\rho_{t}(x_{t})$ $=(1-t)^{-1}\int\nabla\log\rho_{0}(x_{0})\rho(x_{0},x_{1}|x_{t})\mathrm{d}x_{0}\mathrm{d}x_{1}$ (7)
$=t^{-1}\int\nabla\log\rho_{1}(x_{1})\rho(x_{0},x_{1}|x_{t})\mathrm{d}x_{0}\mathrm{d}x_{1},$ (8)
combined with the availability of the target scores, yields the objectives
$\mathcal{L}^{0}_{\mathrm{TSM}}$ $=\mathbb{E}_{x_{0},x_{1},x_{t}}||\nabla_{x}U_{t}(x_{t})+\tfrac{1}{1-t}\nabla\log\rho_{0}(x_{0})||^{2}$ (9)
$\mathcal{L}^{1}_{\mathrm{TSM}}$ $=\mathbb{E}_{x_{0},x_{1},x_{t}}||\nabla_{x}U_{t}(x_{t})+\tfrac{1}{t}\nabla\log\rho_{1}(x_{1})||^{2}.$ (10)
In our experiments we combine these two objectives and work with the following target score matching objective,
$\mathcal{L}_{\mathrm{TSM}}=\mathbb{I}_{t<0.5}\mathcal{L}^{0}_{\mathrm{TSM}}+\mathbb{I}_{t\geq 0.5}\mathcal{L}^{1}_{\mathrm{TSM}}.$ (11)
@@
Both SM objectives require the computation of the spatial gradient $\nabla_{x}U_{t}(x)$ at $(t,x)\sim(\mathcal{U}([0,1]),\rho_{t})$. The temporal derivative of the energy function, $\partial_{t}U_{t}(x)$, can be easily calculated thanks to automatic differentiation at the same points, yielding a Monte Carlo estimate of the TI integral in Equation (2),
$\Delta\hat{F}_{0\rightarrow 1}=\mathbb{E}_{t\sim\mathcal{U}([0,1]),x\sim\rho_{t}}\big{[}\partial_{t}U_{t}(x)\big{]}.$ (12)
In our experiments, at every training step we compute both the spatial and temporal derivatives of $U_{t}$ and use $\nabla U_{t}$ for the SM objective and $\partial_{t}U_{t}$ for a rolling estimate of the free-energy difference.
Solvation free energies
@@
$U_{1}$ $=U_{\mathrm{solvent}}+U_{\mathrm{solute}}+U_{\mathrm{solute-solvent}},$ (14)
where $U_{\mathrm{solvent}},U_{\mathrm{solute}}$ and $U_{\mathrm{solute-solvent}}$ represent the interactions within the solvent, within the solute and cross interactions between the two components, respectively. In Section IV we consider two experimental setups, ($i$) an LJ solute included in an LJ solvent and the inclusion of a rigid solute molecule in a solvent of rigid water molecules. We impose periodic boundary conditions, leading to configuration spaces $\mathbb{T}^{3}$ and $[\mathbb{T}^{3}\times SO(3)]^{N}$, respectively. The solute is rigid in both cases, implying $U_{\mathrm{solute}}\equiv 0$. As described below, most properties of these systems are shared. When discussing the differences between them we refer to the two setups by LJ-experiment and hydration-experiment.
Interactions
We consider system where the components interact via Lennard-Jones (LJ) and Coulomb interactions. A softening of these interactions is necessary as the interpolating samples might bring particles close to each other, causing numerical issues with the unsoftened interaction. We use the LJ interaction with softening parameter[36] given by
where $r$ is the distance between the particles and $(\epsilon,\sigma)$ denote the energy and length scales of the LJ interactions. When $a=0$, no softening is applied, and larger values of $a$ mean larger softening deformation of the original Lennard-Jones interactions. Note that equation (15) corresponds to evaluating the unsoftened Lennard-Jones potential at $\sqrt{r^{2}+a\sigma^{2}}$. Interactions between different species, $A$ and $B$, of particles are computed by the Lorentz-Berthelot rules [37, 38] $(\varepsilon_{AB}=\sqrt{\varepsilon_{A}\varepsilon_{B}},\sigma_{AB}=(\sigma_{A}+\sigma_{B})/2)$.
Coulomb interactions are modelled with the reaction field method [39]
$U_{\text{Coulomb}}(r)$ =$=\begin{cases}k_{C}{q_{1}q_{2}}\left(\frac{1}{r}+k_{\text{rf}}r^{2}-c_{\text{rf}}\right),\quad\text{if}\quad r\leq r_{\text{cutoff}}\\ 0,\quad\text{if}\quad r>r_{\text{cutoff}}\end{cases}$ (16)
$k_{\text{rf}}$ $=\left(\frac{1}{r_{\text{cutoff}}^{3}}\right)\left(\frac{\epsilon_{\text{solvent}}-1}{2\epsilon_{\text{solvent}}+1}\right)$ (17)
$c_{\text{rf}}$ $=\left(\frac{1}{r_{\text{cutoff}}}\right)\left(\frac{3\epsilon_{\text{solvent}}}{2\epsilon_{\text{solvent}}+1}\right),$ (18)
where $r_{\text{cutoff}}$ is the cutoff radius of the Coulomb interaction, $k_{C}$ is Coulomb’s constant, $q_{1},q_{2}$ are the charges of the atoms, $r$ is the distance between them and $\epsilon_{\text{solvent}}$ is the dielectric constant of the solvent. To soften this potential with softening parameter $a$, we define
$U^{\text{soft}}_{\text{Coulomb}}(r,a)=U_{\text{Coulomb}}(\sqrt{r^{2}+a}).$ (19)